OpenStudy (anonymous):

Segment TV has endpoints at (2, 10) and (18, -18). What is the length of the segment? Round to the nearest hundredth. 29.00 32.25 49.07 47.92

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

distance formula

ganeshie8 (ganeshie8):

\(\sqrt{(x2-x1)^2 + (y2-y1)^2}\)

ganeshie8 (ganeshie8):

(2, 10) and (18, -18) x1 y1 x2 y2

ganeshie8 (ganeshie8):

plugin

OpenStudy (anonymous):

So (18 - 2)^2 + (-18 - 10)^2

ganeshie8 (ganeshie8):

sqrt of that

OpenStudy (anonymous):

Do I simplify it, and then sqrt it?

ganeshie8 (ganeshie8):

yes. first simplify, then sqrt

OpenStudy (anonymous):

(16)^2 + (-28)^2 = 256 + 784 = 1040 = 32.2490309932

OpenStudy (anonymous):

So 32.35 ?

OpenStudy (anonymous):

32.25*

ganeshie8 (ganeshie8):

yes! gw

OpenStudy (anonymous):

Thank you! Can you help me with one more? Sorry, I am doing a 60 question review

ganeshie8 (ganeshie8):

ohk.. oly if its not long, else we can do tomoro...

OpenStudy (anonymous):

A (4, 4), B (7, 0), C (11, 3), and D (8, 7) are four points on the coordinate grid. Miranda and Pete joined the points using straight lines to draw a quadrilateral ABCD. Which statement is correct? Pete is correct because all the four sides of the quadrilateral ABCD are equal, therefore it is a rhombus. Miranda is incorrect because she has used incorrect formula to calculate the slope of the lines. Pete is incorrect because he has used incorrect formula to find the distance between the points of the segments. Miranda is correct because the slope of AB is equal to DC and slope of BC is equal to AD, therefore it’s a rhombus.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

after proving its a parallelogram - to prove it is not a rhombus, you need to prove adjacent sides are not equal. but Mirinda didnt do that - So he is incorrect.

ganeshie8 (ganeshie8):

this is the answer :- Pete is correct because all the four sides of the quadrilateral ABCD are equal, therefore it is a rhombus.

OpenStudy (anonymous):

Thank you! That was my last question, you really have helped me out these past few days! I appreciate it so much! Have a good night, if it is night where you are (:

ganeshie8 (ganeshie8):

np :D you really deserve good sleep after so much work... gn :)